OpenStudy (anonymous):

Beautiful people; one of you PLEASE HELP ME! :) Will FAN and GIVE MEDAL! The kitchen in Jerry’s house is a square. The points P(1, 7) and Q(5, 1) represent the opposite vertices of his kitchen in a drawing on a coordinate grid. Which of these ordered pairs could be a third vertex? A. (7, 6) B. (0, -2) C. (6, 6) D.(1, 2) @ganeshie8 @epoweritheta @joshhfers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is his kitchen triangular ?

OpenStudy (anonymous):

Read the first line..

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

it's a square

geerky42 (geerky42):

Plot on graph would help.

OpenStudy (anonymous):

That's all I was given :/

OpenStudy (anonymous):

it (6, 6)

OpenStudy (anonymous):

thats what im sayin

OpenStudy (anonymous):

it's ^*

geerky42 (geerky42):

You guys....

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

How is it (6,6?)

OpenStudy (anonymous):

Yes @geerky42 ?

OpenStudy (anonymous):

The distance between two points P and Q is sqrt((x1-x2)^2 + (y1-y2)^2) : In this case (1-5)^2 + (7-1)^2 = 36+ 16 = 52 (PQ^2) okay?

OpenStudy (anonymous):

Alrighty thank you, I need help on this last one and I'm good... An artist designed a badge for a school club. Figure ABCD on the coordinate grid below shows the shape of the badge.

OpenStudy (anonymous):

i still wasn't done explaining it lol but yah let me see if i can do it and are you flvs student?

OpenStudy (anonymous):

Oh..sorry and no, I take another online server. The badge is enlarged and plotted on the coordinate grid as Figure JKLM with point J located at (-4, -7) and point M located at (-2, -3). Which of these could be the possible coordinates for vertex L? A. (-1, -7) B. (3, -7) C. (0, -3) D. (2, -3)

OpenStudy (anonymous):

i think B but not sure but give me one second

OpenStudy (anonymous):

no no it's not B

OpenStudy (anonymous):

do you have a clue of what it might be? @lilia222

OpenStudy (anonymous):

No, this is why I'm seeking all the help I could get..

OpenStudy (anonymous):

should i tell you how to get it instead of giving you the answer?

OpenStudy (anonymous):

On the bright side, at least we can eliminate choice "B" .. :))

OpenStudy (anonymous):

I rather learn how to do it..

OpenStudy (anonymous):

Well the second enlarged diagram is twice as big as the first one.The length of AD=sqrt(5) and JM=2*sqrt(5). So the sides are in the ratio of 1:2.So ML will be twice as long as DC.

OpenStudy (anonymous):

once you get the answer i would let you know if its correct or not

OpenStudy (anonymous):

can i get a medal for answering the other question?

OpenStudy (anonymous):

Yes! I'm confused how you do the 2nd question though...

OpenStudy (anonymous):

Could you do the steps with me? :)

OpenStudy (anonymous):

okay well it's C and i explained it above you can look back at it and it will probably make more sense since you already got the answer (0, -3)

OpenStudy (anonymous):

Thank you ooohh sooooo sooooooo much! Seriously ! :))

OpenStudy (anonymous):

sh*t i shouldn't have given you the answer but if you want me i can still go over it with you

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

-but you explained it above; I'm going to tag you in another one-just for the 2nd ques.